• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name ejensen

    Log entry time 21:31:05 on July 9,2010

    Entry number 331175

    This entry is a followup to: 330740

    keyword=Wierd timing peaks from earlier

    It looks like the double peak issue in the left arm timing spectrum has disappeared. The first plot is of the s2 timing spectrum from a run taken yesterday morning. This was the last run before entering into restricted access. The second plot is from the first run taken after we got beam back yesterday evening. I realize these plots are not totally convincing, but I have studied this using different cuts and variables and arrive at the same conclusion. Perhaps someone fixed the problem during the access yesterday.

    FIGURE 1

    FIGURE 2